OpenStudy (anonymous):

Solve the system of equations. 9x + y = 2 –4x – y = –17 A. (–3, 29) B. (1, –7) C. (0, 0) D. (4, 1)

OpenStudy (anonymous):

9x ? 2 = 4x + 3 9x - 4x = 3 + 2 5x = 5 x = 5/5 x = 1